# Bit Threads: From Entanglement to Geometric Entropies

### v1

@@ 2.2 Bit threads from CPS
In the previous subsection we saw that how one can find the (d2)(d-2)-form $\mathbf{k}_{\xi}$ (and $\boldsymbol{\mathit{j}}_{\xi}$) from the covariant phase space formalism. Here, we will show how to construct a flow from this form that is a potential bit threads vector field. Both of the above mentioned CPS formalisms are equally valid for this construction as they lead to the same entropy. We will mainly follow the IW prescription. The codimension2-2 current $\mathbf{k}_{\xi}$ is given by
@@
for any diffeomorphism generator $\xi^{\mu}$. From hereafter we will be omitting the suffix ‘IW’ to simplify the notations. In the covariant phase space formalism [15, 23] these codim-2 currents are used to obtain variations of different conserved charges in a black hole background which lead to the renowned first law of black hole thermodynamics (see appendix B.3 for a short review). We can extend this method to obtain the conserved charges by using the (d2)(d-2)-form $\boldsymbol{\mathit{j}}_{\xi}$ (if present) instead of the usual $\mathbf{k}_{\xi}$. We mainly focus on the exact symmetries (like isometry) of the theory. When $\xi^{\mu}$ is a Killing vector (or any exact symmetry generator), the symplectic current $\boldsymbol{\omega}(\delta\Phi,\mathcal{L}_{\xi}\Phi,\Phi)$ vanishes, i.e.,
@@
where we have denoted $\eta$’ as the set of all Killing vectors $\xi$’s. The above equation implies that $\mathbf{k}_{\eta}(\delta\Phi,\Phi)$ is a closed form and so is $\boldsymbol{\mathit{j}}_{\eta}$. Then in the presence of a Riemannian metric $g_{ab}$ ($a,b=1,\cdots,d-1$), we can have a divergenceless vector field $v^{a}$ from both the closed forms $\mathbf{k}_{\eta}(\delta\Phi,\Phi)$ and $\boldsymbol{\mathit{j}}_{\eta}$. We will use the latter one to obtain a vector field using the map:
@@
where, $\star$’ is the Hodge dual defined on a Cauchy surface. Again, we can write [25],
@@
where $\boldsymbol{\epsilon}$ is the volume form of the Cauchy slice. As $\boldsymbol{\mathit{j}}_{\eta}$ is a closed form, it implies that the vector field is indeed divergenceless (i.e., $\nabla_{a}v_{cps}^{a}=0$) everywhere.444 On the uniqueness of $\boldsymbol{v_{cps}^{a}}$ : The fundamental identity (2.2) allows us to define the codim-2 current $\mathbf{k}_{\xi}$ from the symplectic current, but this definition involves some ambiguities that we need to address. The symplectic potential $\mathbf{\Theta}$ itself is not uniquely defined [15]. Adding boundary terms to the Lagrangian changes $\mathbf{\Theta}$ but leaves the symplectic current $\boldsymbol{\omega}=\delta\mathbf{\Theta}$ unchanged. More generally, we can modify $\mathbf{\Theta}\rightarrow\mathbf{\Theta}-\mathrm{d}\mathbf{B}$, which transforms the symplectic current as $\boldsymbol{\omega}\rightarrow\boldsymbol{\omega}+\mathrm{d}\boldsymbol{\omega}_{B}$. These are called “residual ambiguities” and reflect our freedom in choosing boundary terms [26]. However, for exact symmetries (such as Killing symmetries), these ambiguities cancel out. When we transform the symplectic current $\boldsymbol{\omega}\rightarrow\boldsymbol{\omega}+\mathrm{d}\boldsymbol{\omega}_{B}$, the codim-2 current transforms as: $\mathbf{k}_{\xi}\rightarrow\mathbf{k}_{\xi}+\boldsymbol{\omega}_{B}(\delta\Phi,\mathcal{L}_{\xi}\Phi,\Phi)$ For Killing vectors or generators of exact symmetries, the Lie derivative $\mathcal{L}_{\xi}\Phi$ equals the field variation $\delta_{\xi}\Phi$, so the additional term vanishes and $\mathbf{k}_{\xi}$ remains unchanged. This means the vector field $v_{cps}^{a}$ obtained through the CPS construction is uniquely defined (up to the usual degrees of freedom in $\mathbf{k}_{\xi}$) when we work with exact symmetries. Equivalently we can have the relation at the boundary of a Cauchy surface as,
@@
where, $\tilde{\boldsymbol{\epsilon}}$ is volume form on its boundary. We use the (d2)(d-2)-form $\mathbf{k}$ to construct a vector field which will help us to study thermodynamics of a gravitational system. We will discuss this briefly in the section 4.
A central requirement of the bit thread formalism is that $v^{a}$ must be divergenceless. The vector fields derived from the CPS formalism using the codim-2 form $\mathbf{k}_{\xi}$ (or $\boldsymbol{\mathit{j}}_{\xi}$) automatically fulfil this condition. This suggests that imposing the remaining bit thread constraints onto these CPS-derived flows might provide a canonical construction for bit threads. To have a better understanding of the connection between these two formalisms, we must re-examine the properties of the codim-2 form $\mathbf{k}_{\xi}$. Although $\mathbf{k}_{\xi}$ resolves ambiguities present in the symplectic current (as explained in the footnote 4), it can be expressed upto an exact form. This freedom in $\mathbf{k}_{\xi}$ can be utilised to satisfy the properties of the bit threads. So we can write the current form as
@@
The same ensues to the vector field obtained from the above current form:
@@
where, $v^{a}_{\mathrm{d}}$ is the vector field corresponding to the co-exact $1$-form $\star\mathrm{d}\boldsymbol{\mathit{j}}_{B_{\xi}}$ and is divergence free. Now, if the given manifold (or the domain of our interest) is simply connected, $v^{a}_{\mathrm{d}}$ can be expressed as a gradient of some function $\varphi(x)$. So, the term $v_{\mathrm{d}}^{a}=\nabla^{a}\varphi$ is like a $U(1)$ gauge freedom [^5] We thank Mohd Ali for emphasising the importance of this term and for valuable discussions.. This freedom helps us to satisfy the boundary conditions for bit threads in some cases. We are mainly interested in the regions of the spacetime manifold where the domain is simply connected. Then, we can say that $v^{a}$ and $v^{a}_{cps}$ are from the same equivalence class :
@@
Evidently, $v^{a}$ is a divergenceless vector field and we will show that this becomes a valid bit thread configuration once we impose the conditions (1.2b). If the manifold has non-trivial topologies (when the manifold is not simply connected), the bit threads field $v^{a}$ and the vector derived from CPS $(v^{a}_{cps})$ differ by a divergenceless vector field $v^{a}_{\mathrm{d}}$. Therefore, in general the equivalence class of $v^{a}_{cps}$ can be written as: $v^{a}_{gen}]=\{v^{a}_{cps}\,|\,\,v^{a}_{cps}\sim v^{a}_{cps}+v^{a}_{\mathrm{d}}\,\}$ with $\nabla_{a}v^{a}_{\mathrm{d}}=0$.
The bit threads vector field that we obtained using the (d2)(d-2)-form $\boldsymbol{\mathit{j}}$ arising from the CPS formalism is called ‘canonical bit thread’. Although this formalism does not require any notion of metric until we switch to the notion of a vector: the map (2.13) precisely use the metric of a Cauchy surface to construct the desired $v_{cps}^{a}$ from the differential form $\boldsymbol{\mathit{j}}$. And it is essential to invoke the non-trivial $v^{a}_{\mathrm{d}}$ to satisfy all the conditions of bit threads. For examples we will look into AdS3 spacetime in two different coordinate systems.
@@
where, $\xi^{\mu}$ is a Killing vector, $h^{\mu\nu}=g^{\mu\sigma}g^{\nu\gamma}\,\delta g_{\sigma\gamma}$ is the metric perturbation and $h\equiv h^{\mu}_{\mu}=g_{\mu\nu}h^{\mu\nu}$ is its trace. These indices run from 0 to d1d-1 for a $d$ dimensional spacetime. The codimension-2 form is then given by,
@@
where, $P$ and $Q$ are some parameters of the metric. For the present case (2.19), the only constant we have is the AdS radius $L$. So, the parametric variation looks like $\hat{\delta}g_{\mu\nu}(L)=\dfrac{\partial g_{\mu\nu}}{\partial L}\hat{\delta}L$. Using this parametric variation, we find the only non-zero component of the form $\mathbf{k}_{\xi}$ :
@@
We can also find another codim-2 current $\boldsymbol{\mathit{j}}_{\xi}$ such that $\mathbf{k}_{\xi}=-\hat{\delta}\boldsymbol{\mathit{j}}_{\xi}$ given by
@@
Now using the relation (2.13), we can find a divergenceless vector field $v_{cps}^{a}$ to be [^6] We thank Manish Ramchander for pointing out a mistake in this expression at the earlier stage of the work.,
@@
where, $\mathfrak{B}$ is an overall constant and is a function of $L$. We have kept the factor $1/4G_{N}$ outside the constant to keep track of the constraints (1.2b). From this expression it is clear that $v_{cps}^{a}$ is divergenceless. But it fails to be proportional to the $n^{a}$ on the RT surface. Recall the RT surface in Poincaré coordinates is given by $x^{2}+z^{2}=R^{2}$, $R$ being the radius of the RT surface. Then the unit normal is $n^{a}=(n^{z},n^{x})=\frac{z_{m}}{R}(z_{m},x_{m})$, where $z_{m}$ and $x_{m}$ are the coordinates on the minimal surface. One can easily check that $v_{cps}^{a}\neq n^{a}/4G_{N}$ on the RT surface. So $v^{a}_{cps}$ alone can not reproduce the HEE. But we can use the gauge freedom $v^{a}_{\mathrm{d}}=\nabla^{a}\varphi$ associated with $v^{a}_{cps}$ to make it a bit thread vector field. To find $v_{\mathrm{d}}^{a}$, we first solve $\nabla_{a}\nabla^{a}\varphi=0$ in the Poincaré background, and then substitute $v^{a}_{\mathrm{d}}$ with the gradient of $\varphi$. Solving of the Laplace’s equation with proper boundary conditions gives us the harmonic function
$\varphi=\dfrac{l^{2}_{\rm{AdS}}}{4G_{N}R}\left(z-\dfrac{\mathfrak{B}z^{3}}{6}+\mathfrak{B}\dfrac{zx^{2}}{2}+\dfrac{x\sqrt{R^{2}-z^{2}}}{z^{2}}\left(z-\mathfrak{B}z^{3}\right)\right).$ (2.23)
Taking gradient of the above function and putting together with (2.22), we find the total vector field,
@@
which reduces to $n^{a}$ on the RT surface and also satisfies the property $|v|\leq 1/4G_{N}$. So this is a flow which correctly reproduces the HEE of a single interval in Poincaré AdS. To interpret the different parts of $v^{a}$, we can say that the $v_{cps}^{a}$ store all the information regarding the conserved charges of the theory as it is derived from the symplectic current, and $\nabla^{a}\varphi$ helps to satisfy the constraints of bit threads. From this example, it is clear that $v^{a}$ and $v^{a}_{cps}$ are from the same equivalence class, i.e., $v^{a}\sim v^{a}_{cps}$.
A similar treatment which we used in the Poincaré background can be done for other backgrounds as well. As an example we discuss the case of D11-brane in the appendix C. There can be other ways to find the $v_{\mathrm{d}}^{a}$ in different spacetimes which are more convenient to use than the method presented here. In case of Poincaré $\rm{AdS}$, we can directly obtain a flow solely using the methods of characteristics starting from the divergenceless condition. We briefly discuss it in appendix D.
@@
In the previous subsection, we have seen that it was required to use the freedom presented in the definition of $v_{cps}^{a}$ to satisfy all the conditions to be a flow even in a simple geometry like Poincaré AdS. So, a natural question arises: Can we find geometries where $v_{cps}^{a}$ is sufficient to be a flow without the aid of its gauge freedoms? To answer this question, we need to go to the reference frame of an observer whose causal boundary is bounded by the RT surface. Then in the natural coordinates of that observer, we can obtain a valid bit thread vector field from the CPS formalism. One can always map a spherical region of the boundary CFT to this kind of wedge in the hyperbolic foliation of the dual AdS spacetime [27]. For example, consider an accelerated observer in $\rm{AdS}_{3}$ spacetime. The reference frame of the observer is the Rindler-$\rm{AdS}_{3}$ geometry,
@@
This form is defined in the whole spacetime. By projecting it on a Cauchy surface, and using the map (2.13), we find
@@
By definition, the thread lines start at the boundary subregion (here the whole boundary) and end at the complement of that subregion and horizon or any other codimension-2 surface present in the spacetime. But the above vector $v^{\rho}$ is directed towards the boundary as that is the direction in which $\rho$ increases. So the actual vector field will be
@@
The above vector field is divergenceless by definition and satisfies the necessary conditions for a valid bit thread configuration: $|v|\leq 1/4G_{N},$ and $v^{a}=-n^{a}/4G_{N}$ on the horizon [^7] Here we are having an extra minus sign because the unit normal is pointing outward at the horizon whereas the direction of the bit thread is towards the horizon. of Rindler space (i.e., the RT surface). These conditions ensure consistency with the bit thread formalism (see equation (A.1)), allowing us to conclude that this vector field constitutes a physically admissible bit thread flow.
@@
Here we will show how to obtain a bit thread vector field in Poincaré coordinates from the vector (2.27) by using the coordinate transformations presented in appendix E. First, we have to make sure that the metrics presented earlier and in the appendix are same to use the transformations. Comparing the metrics (2.25) and (E.8), we set the acceleration parameter $\mathtt{a}=1$ in (2.25) which yields
@@
$v^{a}=\left(-v^{\rho},v^{u}\right)=-\dfrac{1}{4G_{N}}\left(\dfrac{L}{\rho}\sqrt{\dfrac{\rho^{2}}{L^{2}}-1},0\right)=\dfrac{1}{4G_{N}}\left(\dfrac{L}{\rho}\sqrt{\dfrac{\tilde{\rho}^{2}}{L}},0\right).$ (2.28)
@@
$\vec{v}=v^{a}\partial_{a}=\underbrace{\left(-v^{\rho}\partial_{\rho},v^{u}\partial_{u}\right)}_{Rindler\,AdS}=\underbrace{\left(v^{z}\partial_{z},v^{x}\partial_{x}\right)}_{Poincar\acute{e}\,AdS}$
@@
where, $v^{a}_{geodesic}$ is the symmetric flow constructed using geodesics in [18]. This vector field is divergenceless, obeys the norm bound $|v|\leq\frac{1}{4G_{N}L}$, and at the RT surface reduces to $v^{a}|_{RT}=\frac{n^{a}}{4G_{N}L}$. Evidently, it is a valid thread configuration. It is surprising that only using the coordinate transformations we found a thread configuration in Poincaré background without any prior knowledge of the RT surface, whereas in constructing the symmetric flow from geodesics the knowledge of the RT surface were required as shown in [18].
Figure 1: Flow lines (black curves with arrows) for bit threads obtained from transforming a valid thread configuration in AdS-Rindler to Poincaré. The gray horizontal line denotes the boundary and the red semicircle is the RT surface corresponding to the boundary subregion (shown in blue) $A\in[-2,2]$.
So far, we have found a thread configuration for a fixed length interval $L$, but in general there can be any size of interval in the boundary CFT. In the bulk theory we have corresponding RT surfaces of different radii. These RT surfaces can be related via the isometries of the gravity theory. As discussed in appendix E, one such simple isometry is boost (E.13) in the embedding plane $(Y_{-1},Y_{2})$. We can find the corresponding vector field in this case as well by proceeding the similar way as above. In the boosted case the expression (2.30) becomes
@@
And using the coordinate relations (E.17), we get
@@
Now, if we identify $R=e^{-\beta}L$ to be the radius of the entangling surface in the boundary theory, we end up with a vector field proportional to $v^{a}_{geodesic}$:
@@
The flow lines of this vector field are shown in fig. 1. The HEE corresponding to a boundary subregion $A$ in terms of the above thread configuration can be written as,

### v2

@@ 2.2 Bit threads from CPS
In the previous subsection, we saw how one can find the (d2)(d-2)-form $\mathbf{k}_{\xi}$ (and $\boldsymbol{\mathit{j}}_{\xi}$) from the covariant phase space formalism. Here, we will show how to construct a flow from this form, which is a potential bit thread vector field. Both of the above-mentioned CPS formalisms are equally valid for this construction, as they lead to the same entropy. We will mainly follow the IW prescription. The codimension2-2 current $\mathbf{k}_{\xi}$ is given by
@@
for any diffeomorphism generator $\xi^{\mu}$. From hereafter, we will be omitting the suffix ‘IW’ to simplify the notations. In the covariant phase space formalism [39, 29], these codim-2 currents are used to obtain variations of different conserved charges in a black hole background, which lead to the renowned first law of black hole thermodynamics (see appendix B.3 for a short review). We can extend this method to obtain the conserved charges by using the (d2)(d-2)-form $\boldsymbol{\mathit{j}}_{\xi}$ (if present) instead of the usual $\mathbf{k}_{\xi}$. We mainly focus on the exact symmetries (such as isometry) of the theory. When $\xi^{\mu}$ is a Killing vector (or any exact symmetry generator), the symplectic current $\boldsymbol{\omega}(\delta\Phi,\mathcal{L}_{\xi}\Phi,\Phi)$ vanishes, i.e.,
@@
where we have denoted $\eta$’ as the set of all Killing vectors. The above equation implies that $\mathbf{k}_{\eta}(\delta\Phi,\Phi)$ is a closed form and so is $\boldsymbol{\mathit{j}}_{\eta}$. Then in the presence of a Riemannian metric $g_{ab}$ ($a,b=1,\cdots,d-1$), we can have a divergenceless vector field $v^{a}$ from either of the closed forms $\mathbf{k}_{\eta}(\delta\Phi,\Phi)$ or $\boldsymbol{\mathit{j}}_{\eta}$. We will use the latter one to obtain a vector field using the map: [^6] On the uniqueness of $\boldsymbol{v_{cps}^{a}}$ : The fundamental identity (2.2) allows us to define the codim-2 current $\mathbf{k}_{\xi}$ from the symplectic current, but this definition involves some ambiguities that we need to address. The symplectic potential $\mathbf{\Theta}$ itself is not uniquely defined [39]. Adding boundary terms to the Lagrangian changes $\mathbf{\Theta}$ but leaves the symplectic current $\boldsymbol{\omega}=\delta\mathbf{\Theta}$ unchanged. More generally, we can modify $\mathbf{\Theta}\rightarrow\mathbf{\Theta}-\mathrm{d}\mathbf{B}$, which transforms the symplectic current as $\boldsymbol{\omega}\rightarrow\boldsymbol{\omega}+\mathrm{d}\boldsymbol{\omega}_{B}$. These are called “residual ambiguities” and reflect our freedom in choosing boundary terms [18]. However, for exact symmetries (such as Killing symmetries), these ambiguities cancel out. When we transform the symplectic current $\boldsymbol{\omega}\rightarrow\boldsymbol{\omega}+\mathrm{d}\boldsymbol{\omega}_{B}$, the codim-2 current transforms as: $\mathbf{k}_{\xi}\rightarrow\mathbf{k}_{\xi}+\boldsymbol{\omega}_{B}(\delta\Phi,\mathcal{L}_{\xi}\Phi,\Phi)$ For Killing vectors or generators of exact symmetries, the Lie derivative $\mathcal{L}_{\xi}\Phi$ equals the field variation $\delta_{\xi}\Phi$, so the additional term vanishes and $\mathbf{k}_{\xi}$ remains unchanged. This means the vector field $v_{cps}^{a}$ obtained through the CPS construction is uniquely defined (up to the usual degrees of freedom in $\mathbf{k}_{\xi}$) when we work with exact symmetries.
@@
where, $\star$’ is the Hodge dual defined on a Cauchy surface. [^7] We will mostly work in the spacetimes that are asymptotically AdS. AdS spacetime is not globally hyperbolic and therefore does not admit a Cauchy surface, which is required to have a well-defined initial value problem and hence the canonical phase space. However, a well-defined phase space can be constructed in AdS by choosing suitable boundary conditions [38, 56]. So, we will use the term ‘Cauchy’ surface to refer to the spacelike hypersurfaces with proper boundary conditions in asymptotic AdS spacetimes. Again, we can write [57],
@@
where $\boldsymbol{\epsilon}$ is the volume form of the Cauchy slice. As $\boldsymbol{\mathit{j}}_{\eta}$ is a closed form, it implies that the vector field is indeed divergenceless (i.e., $\nabla_{a}v_{cps}^{a}=0$) everywhere. Equivalently, we can have the relation at the boundary of a Cauchy surface as
@@
where, $\tilde{\boldsymbol{\epsilon}}$ is volume form on its boundary. We will use the (d2)(d-2)-form $\mathbf{k}$ to construct a vector field to study the thermodynamics of a gravitational system. We will discuss this briefly in the section 4.
A central requirement of the bit thread formalism is that $v^{a}$ must be divergenceless. The vector fields derived from the CPS formalism using the codim-2 form $\mathbf{k}_{\xi}$ (or $\boldsymbol{\mathit{j}}_{\xi}$) automatically fulfil this condition, which suggests that imposing the remaining bit thread constraints on these CPS-derived flows might provide a canonical construction for bit threads. To have a better understanding of the connection between these two formalisms, we must re-examine the properties of the codim-2 form $\mathbf{k}_{\xi}$. Although $\mathbf{k}_{\xi}$ resolves ambiguities present in the symplectic current (as explained in the footnote 6), it can be expressed up to an exact form. This freedom in $\mathbf{k}_{\xi}$ can be utilised to satisfy the properties of the bit threads. So we can write the current form as
@@
The same ensues for the vector field obtained from the above current form:
@@
where, $v^{a}_{\mathrm{d}}$ is the vector field corresponding to the co-exact $1$-form $\star\mathrm{d}\boldsymbol{\mathit{j}}_{B_{\xi}}$ and is divergence free. Now, if the given manifold (or the domain of interest) is simply connected, $v^{a}_{\mathrm{d}}$ can be expressed as a gradient of some function $\varphi(x)$. So, the term $v_{\mathrm{d}}^{a}=\nabla^{a}\varphi$ is like a $U(1)$ gauge freedom, [^8] We thank Mohd Ali for emphasising the importance of this term and for valuable discussions. which will help us to satisfy the boundary conditions for bit threads in some cases. We are mainly interested in the regions of the spacetime manifold where the domain is simply connected. Then, we can say that $v^{a}$ and $v^{a}_{cps}$ are from the same equivalence class:
@@
Evidently, $v^{a}$ is a divergenceless vector field, and we will show that this becomes a valid bit thread configuration once we impose the conditions (1.2b). If the manifold has non-trivial topologies (when the manifold is not simply connected), the bit threads field $v^{a}$ and the vector derived from CPS $(v^{a}_{cps})$ differ by a divergenceless vector field $v^{a}_{\mathrm{d}}$. Therefore, in general the equivalence class of $v^{a}_{cps}$ can be written as: $v^{a}_{gen}]=\{v^{a}_{cps}\,|\,\,v^{a}_{cps}\sim v^{a}_{cps}+v^{a}_{\mathrm{d}}\,\}$ with $\nabla_{a}v^{a}_{\mathrm{d}}=0$.
The bit thread vector field that we have obtained using the (d2)(d-2)-form $\boldsymbol{\mathit{j}}$ arising from the CPS formalism is called the ‘canonical bit threads’. Although this formalism does not require any metric until we switch to the notion of a vector, the map (2.13) precisely uses the metric of a Cauchy surface to construct the desired $v_{cps}^{a}$ from the differential form $\boldsymbol{\mathit{j}}$. And it is essential to invoke the non-trivial $v^{a}_{\mathrm{d}}$ to satisfy all the conditions of bit threads. For example, we will look into AdS3 spacetime in two different coordinate systems.
@@
where, $\xi^{\mu}$ is a Killing vector, $h^{\mu\nu}=g^{\mu\sigma}g^{\nu\gamma}\,\delta g_{\sigma\gamma}$ is the metric perturbation and $h\equiv h^{\mu}_{\mu}=g_{\mu\nu}h^{\mu\nu}$ is its trace. These indices run from 0 to d1d-1 for a $d$-dimensional spacetime. The codimension-2 form is then given by,
@@
where $P$ and $Q$ are some parameters of the metric. For the present case (2.19), the only constant we have is the AdS radius $L$. So, the parametric variation looks like $\hat{\delta}g_{\mu\nu}(L)=\dfrac{\partial g_{\mu\nu}}{\partial L}\hat{\delta}L$. Using this parametric variation, we find the only non-zero component of the form $\mathbf{k}_{\xi}$ :
@@
We can also find another codim-2 current $\boldsymbol{\mathit{j}}_{\xi}$ such that $\mathbf{k}_{\xi}=\hat{\delta}\boldsymbol{\mathit{j}}_{\xi}$ given by
@@
Now using the relation (2.13), we obtain a divergenceless vector field $v_{cps}^{a}$ given as [^9] We thank Manish Ramchander for pointing out a mistake in this expression at the earlier stage of the work.
@@
where, $\mathfrak{B}$ is an overall constant and is a function of $L$. We have kept the factor $1/4G_{N}$ separate to keep track of the constraints (1.2b). From this expression, it is clear that $v_{cps}^{a}$ is divergenceless. But it fails to be proportional to the unit normal $n^{a}$ on the RT surface. Recall the RT surface in Poincaré coordinates is given by $x^{2}+z^{2}=R^{2}$, $R$ being the radius of the RT surface. Then the unit normal is $n^{a}=(n^{z},n^{x})=\frac{z_{m}}{R}(z_{m},x_{m})$, where $z_{m}$ and $x_{m}$ are the coordinates on the minimal surface. One can easily check that $v_{cps}^{a}\neq n^{a}/4G_{N}$ on the RT surface. So, $v^{a}_{cps}$ alone cannot reproduce the HEE. However, we can utilise the gauge freedom $v^{a}_{\mathrm{d}}=\nabla^{a}\varphi$ associated with $v^{a}_{cps}$ to make it a bit thread vector field. To find $v_{\mathrm{d}}^{a}$, we first solve $\nabla_{a}\nabla^{a}\varphi=0$ in the Poincaré background, and then substitute $v^{a}_{\mathrm{d}}$ with the gradient of $\varphi$. Solving Laplace’s equation with proper boundary conditions gives us the harmonic function
$\varphi=\dfrac{L^{2}}{4G_{N}R}\left(z-\dfrac{\mathfrak{B}z^{3}}{6}+\mathfrak{B}\dfrac{zx^{2}}{2}+\dfrac{x\sqrt{R^{2}-z^{2}}}{z^{2}}\left(z-\mathfrak{B}z^{3}\right)\right).$ (2.23)
Taking the gradient of the above function and putting together with (2.22), we find the total vector field,
@@
which reduces to $n^{a}$ on the RT surface and also satisfies the property $|v|\leq 1/4G_{N}$. So this is a flow which correctly reproduces the HEE of a single interval on the boundary of Poincaré AdS. To interpret the different parts of $v^{a}$, we can say that the $v_{cps}^{a}$ store all the information regarding the conserved charges of the theory as it is derived from the symplectic current, and $\nabla^{a}\varphi$ helps to satisfy the constraints of bit threads. From this example, it is clear that $v^{a}$ and $v^{a}_{cps}$ are from the same equivalence class, i.e., $v^{a}\sim v^{a}_{cps}$.
A similar treatment, which we used in the Poincaré background, can also be done for other backgrounds. As an example, we discuss the case of D11-brane in appendix C. There may be alternative methods to find $v_{\mathrm{d}}^{a}$ in different spacetimes that are more convenient to use than the one presented here. In case of Poincaré $\rm{AdS}$, we can directly obtain a flow solely using the methods of characteristics starting from the divergenceless condition. We briefly discuss it in appendix D.
@@
In the previous subsection, we have seen that it was required to use the freedom presented in the definition of $v_{cps}^{a}$ to satisfy all the conditions to be a flow even in a simple geometry like Poincaré AdS. So, a natural question arises: Can we find geometries where $v_{cps}^{a}$ is sufficient to be a flow without the aid of its gauge freedoms? To answer this question, we need to go to the reference frame of an observer whose causal boundary is bounded by the RT surface. Then, in the natural coordinates of that observer, we can obtain a valid bit thread vector field from the CPS formalism. One can always map a spherical region of the boundary CFT to this kind of wedge in the hyperbolic foliation of the dual AdS spacetime [17]. For example, consider an accelerated observer in $\rm{AdS}_{3}$ spacetime. The reference frame of the observer is the Rindler-$\rm{AdS}_{3}$ geometry,
@@
This form is defined in the whole spacetime. By projecting it on a Cauchy surface and using the map (2.13), we find
@@
By definition, the thread lines start at the boundary subregion (here, the whole boundary) and end at the complement of that subregion and horizon (or any other codim-2 surface present in the spacetime). But the above vector $v^{\rho}$ is directed towards the boundary, as that is the direction in which $\rho$ increases. So the actual vector field will be
@@
The above vector field is divergenceless by definition and satisfies the necessary conditions for a valid bit thread configuration: $|v|\leq 1/4G_{N},$ and $v^{a}=-n^{a}/4G_{N}$ on the horizon [^10] Here we have an extra minus sign because the unit normal is pointing outward at the horizon whereas the direction of the bit thread is towards the horizon. of Rindler space (i.e., the RT surface). These conditions ensure consistency with the bit thread formalism (see equation (A.1)), allowing us to conclude that this vector field constitutes a physically admissible bit thread flow.
@@
Here we will show how to obtain a bit thread vector field in Poincaré coordinates from the vector (2.27) by using the coordinate transformations presented in appendix E. First, we must ensure that the metrics presented earlier and in the appendix are consistent to use the transformations. Comparing the metrics (2.25) and (E.8), we set the acceleration parameter $\mathtt{a}=1$ in (2.25) which yields
@@
$v^{a}=\left(-v^{\rho},v^{u}\right)=-\dfrac{1}{4G_{N}}\left(\dfrac{L}{\rho}\sqrt{\dfrac{\rho^{2}}{L^{2}}-1},0\right)=-\dfrac{1}{4G_{N}}\left(\dfrac{L}{\rho}\sqrt{\dfrac{\tilde{\rho}^{2}}{L}},0\right).$ (2.28)
@@
$\vec{v}=v^{a}\partial_{a}=\underbrace{\left(-v^{\rho}\partial_{\rho},v^{u}\partial_{u}\right)}_{Rindler\,AdS}=\underbrace{\left(v^{z}\partial_{z},v^{x}\partial_{x}\right)}_{Poincar\acute{e}\,AdS}.$
@@
where, $v^{a}_{geodesic}$ is the symmetric flow constructed using geodesics in [3]. This vector field is divergenceless, obeys the norm bound $|v|\leq\frac{1}{4G_{N}L}$, and at the RT surface reduces to $v^{a}|_{RT}=\frac{n^{a}}{4G_{N}L}$. Evidently, it is a valid thread configuration. It is surprising that only using CPS formalism and the coordinate transformations, we found a thread configuration in the Poincaré background without any prior knowledge of the RT surface, whereas in constructing the symmetric flow from geodesics, the knowledge of the RT surface was required as shown in [3].
Figure 1: Flow lines (black curves with arrows) for bit threads obtained from transforming a valid thread configuration in AdS-Rindler to Poincaré. The grey horizontal line denotes the boundary, and the red semicircle is the RT surface corresponding to the boundary subregion (shown in blue) $A\in[-2,2]$.
So far, we have found a thread configuration for a fixed-length interval $L$, but in general, there may be any size of interval in the boundary CFT. In the bulk theory, we have corresponding RT surfaces of different radii. These RT surfaces are related via the isometries of the gravity theory. As discussed in appendix E, one such simple isometry is boost (E.13) in the embedding plane $(Y_{-1},Y_{2})$. We can find the vector field corresponding to one of these intervals by proceeding in a similar manner as above. In the boosted case, the expression (2.30) becomes
@@
And incorporating the coordinate relations (E.17), we get
@@
Now, if we identify $R=e^{-\beta}L$ to be the radius of the entangling surface in the boundary theory (or radius of the RT surface in the bulk), we end up with a vector field proportional to $v^{a}_{geodesic}$:
@@
The flow lines of this vector field are shown in figure 1. The HEE corresponding to a boundary subregion $A$ in terms of the above thread configuration can be written as,